OpenStudy (anonymous):

For the given equation determine the slope and y-intercept of the line algebraically.

OpenStudy (anonymous):

-4x = 9y - 5

OpenStudy (anonymous):

You have to put this in slope-intercept form. -9y = 4x - 5 9y = -4x + 5 y = -4/9x + 5/9 y = mx + b That is the slope intercept form, where m is the slope, and b is the y-intercept. So, -4/9 is the slope, and 5/9 is the y-intercept. :)

OpenStudy (anonymous):

Thanks! Appreciate it.

OpenStudy (anonymous):

No problem! Thanks for the medal. :)